OpenStudy (anonymous):

use a finite sum to estimate the average value of f on the given interval by partitioning the interval into four subintervals of equal length and evaluating the f at the subinterval midpoints. f(x)=x^4 on [0,2]

OpenStudy (anonymous):

Midpoint formula I see, so let's partition 0 .5 1 1.5 2 Since this is midpoint formula,we want to evaluate number in middle 0.25 0.75 1.25 1.75 0.5((0.25)^2+(0.75)^2 +(1.25)^2 +(1.75)^2)

OpenStudy (anonymous):

This get evaluated to be 2.625

OpenStudy (anonymous):

Since this asked for average we multiply by 1/range = 1/2

OpenStudy (anonymous):

i received an answer of 2.63 but wasnt sure this was right. so do i just divide that by 2??

OpenStudy (anonymous):

Yes, divide 2.63 by 2 to get aveage

OpenStudy (anonymous):

\[\text{Avarage}=\frac{1}{b-a}\int _a^bf(x)dx\]

OpenStudy (anonymous):

imranmeah91 do you mind taking a look at mine when u r done?

OpenStudy (zarkon):

3.03515625

OpenStudy (zarkon):

the average vlaue of a function given \[a<x_1<x_2<\cdots<x_n<b\] is \[\frac{\displaystyle\sum_{i=1}^{n}f(x_i)}{n}\]